# Upper bounds on the genus of Albanese fibrations

### v1

@@ Abstract.
Let $S$ be a minimal irregular surface of general type, whose Albanese map induces a hyperelliptic fibration $f:\,S\to B$ of genus $g$. We prove a quadratic upper bound on the genus $g$, i.e., $g\leq h\big{(}\chi(\mathcal{O}_{S})\big{)}$, where $h$ is a quadratic function. We also construct examples showing that the quadratic upper bounds can not be improved to the linear ones. In the special case when $p_{g}(S)=q(S)=1$, we show that $g\leq 14$.
@@
This work is supported by National Natural Science Foundation of China, Fundamental Research Funds for central Universities, and Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014)
@@ 1. Introduction
We work over the complex number throughout this paper. Let $S$ be a minimal irregular surface of general type, and $a:\,S\to\mathrm{Alb}(S)$ its Albanese map. We are interested in the case when the image $a(S)$ is a curve. In this case, the Albanese map induces a fibration, which we call the Albanese fibration of $S$:
@@
In fact, by the universal property of the Albanese map, $B\cong a(S)$, and under this isomorphism the above fibration $f$ is nothing but the Albanese map of $S$. Let $g$ be the genus of a general fiber of $f$. A natural problem is to study the behavior of the genus $g$.
@@
By [Cat00], it is known that the genus $g$ of the Albanese fibration of $S$ is a differential invariant, and hence is also a deformation invariant. Fixing $\chi(\mathcal{O}_{S})$ and $K_{S}^{2}$, there are only finitely many deformation equivalence classes of such surfaces, cf. [BHPV04, § VII]. Hence there must be an upper bound on $g$ depending on $\chi(\mathcal{O}_{S})$ or $K_{S}^{2}$. However, explicit upper bounds are only known under some extra assumptions.
@@
    Suppose that $g(B)=1$ and $K_{S}^{2}=4\chi(\mathcal{O}_{S})$. Ishida [Ish05] proved that $g\leq 2\chi(\mathcal{O}_{S})+2$ if moreover $f$ is hyperelliptic.
@@
Let $S$ be a minimal irregular surface of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$, and $f:\,S\to B$ its Albanese fibration whose general fiber is of genus $g$. Suppose that $f$ is hyperelliptic. Then
$g\leq h\big{(}\chi(\mathcal{O}_{S})\big{)},\qquad\text{where $h$ is some quadratic function.}$.
The above quadratic function can be explicitly written out, see Theorem 4.1. We will also construct in Section 5 examples of minimal irregular surfaces of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$ admitting hyperelliptic Albanese fibrations, whose general fiber is of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$. This is different from the phenomenons when $K_{S}^{2}\leq 4\chi(\mathcal{O}_{S})$, where the known upper bounds are all linear in $\chi(\mathcal{O}_{S})$ as mentioned above.
@@
As mentioned above, the genus $g$ is a deformation invariant when $f$ is the Albanese fibration. Hence one may expect an upper bound on $g$ using invariants of $S$. In general, there is no such upper bound on $g$ if $f$ is not the Albanese fibration. In fact, Penegini-Polizzi ([PP17]) constructed a minimal surface $S$ with $p_{g}(S)=q(S)=2,K_{S}^{2}=4$ (whose Albanese map is generically finite), on which there are fibrations $f_{k}:\,S\to B$, such that the genera of $f_{k}$’s can be arbitrarily large. Nevertheless, in the case when $f:\,S\to B$ is hyperelliptic, the quadratic upper bound proved in Theorem 1.1 holds true even if $f$ is not the Albanese fibration, and hence our results can be applied in a more general situation. See Theorem 4.1 for more details.
Our another interest is the classification of minimal irregular surface of general type with $\chi(\mathcal{O}_{S})=1$. Although the value $\chi(\mathcal{O}_{S})$ is the minimal possibility among surfaces of general type, the classification of such surfaces is widely open. Beauville proved in an appendix to [Deb82] that $p_{g}(S)=q(S)\leq 4$ for such surfaces, and the equality holds if and only if $S$ is isomorphic to a product of two curves of genus $2$. We are more interested in the case when the Albanese map of $S$ induces a fibration. In this case, $p_{g}(S)=q(S)\leq 3$ by Beauville’s result above. In a series of works [HP02, Pen11, Pir02, Zuc03], a full classification has been obtained when $p_{g}(S)=q(S)=3$ or $p_{g}(S)=q(S)=2$. However, the case when $p_{g}(S)=q(S)=1$ seems to be much more mysterious. An explicit upper bound on the genus $g$ of the Albanese fibration is still unknown, although theoretically such an upper bound exists since there is at most finitely many deformation equivalence classes of such surfaces. We prove an explicit upper bound on the genus if the Albanese fibration is hyperelliptic.
@@
Let $S$ be a minimal irregular surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ is hyperelliptic of genus $g$. Then $g\leq 14$. More precisely, it holds $g\leq 10$ except for the following two possible cases:     (i) $K^{2}=8,~{}g=14$;     (ii) $K^{2}=8,~{}g=11$.
The known examples of surfaces with $p_{g}(S)=q(S)=1$ are mostly restricted to the region $g\leq K_{S}^{2}$. The first example with $g>K_{S}^{2}$ was constructed in [FP15], in which $g=7$ and $K_{S}^{2}=6$. More recently, An example with $g=19$ and $K_{S}^{2}=9$ was constructed in [CKY17]. By our result above, the Albanese fibration of the example in [CKY17] must be non-hyperelliptic. The paper is organized as follows. In Section 2, we mainly review some basic facts about surface fibrations and do some technical preparations. In Section 3 and Section 4, we prove Theorem 1.2 and Theorem 1.1 respectively. Finally in Section 5, we construct examples of irregular surfaces of general type with hyperelliptic fibrations of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$.
@@ 2. Preliminaries
In this section, we mainly review some basic facts and fix the notations. In Section 2.1, we recall some general facts about the surface fibrations, and refer to [BHPV04] for more details; and then in Section 2.2 we restrict ourselves to the theory on the hyperelliptic fibrations, which are mainly due to Xiao [Xia91, Xia92].
@@ 2.1. The surface fibrations
Let $f:S\rightarrow B$ be a fibration of curves of genus $g\geq 2$, i.e. $f$ is a proper morphism from a smooth projective surface $S$ onto a smooth projective curve $B$ with connected fibers over the complex number, and the general fiber is a smooth projective curve of genus $g$. The fibration $f$ is called relatively minimal if there is no $(-1)$-curve (i.e. a smooth rational curve with self-intersection $-1$) contained in the fiber of $f$. It is called hyperelliptic if its general fiber is hyperelliptic, smooth if all its fibers are smooth, isotrivial if all its smooth fibers are isomorphic to each other, and locally trivial if it is both smooth and isotrivial. Let $\omega_{S}$ (resp. $K_{S}$) be the canonical sheaf (resp. canonical divisor) of $S$, and let $\omega_{S/B}=\omega_{S}\otimes f^{*}\omega_{B}^{\vee}$ (resp. $K_{f}=K_{S/B}=K_{S}-f^{*}K_{B}$) the relative canonical sheaf (resp. the relative canonical divisor) of $f$. Put $b:=g(B)$, $p_{g}:=h^{0}(S,\omega_{S})$, $q:=h^{1}(S,\omega_{S})$, $\chi=\chi(\mathcal{O}_{S}):=p_{g}-q+1$, and let $\chi_{top}(S)$ be the topological Euler characteristic of $S$. The basic invariants of $f$ are:
@@
    $e_{f}=\sum e_{F}$, where $e_{F}:=\chi_{\rm top}(F_{red})-(2-2g)$ for any fiber $F$ and $F_{red}$ is the reduced part of $F$. Moreover, $e_{F}\geq 0$, and the equality holds if and only if $F$ is smooth. Hence $e_{f}\geq 0$, and $e_{f}=0$ iff $f$ is smooth.
@@ 2.2. Invariants of hyperelliptic fibrations
In this subsection, we restrict to the hyperelliptic fibrations. Let $f:S\rightarrow B$ be a relatively minimal hyperelliptic fibration. The relative canonical map of $f$ is generically of degree 2. Hence it determines an involution $\sigma$ on $S$ whose restriction on a general fiber $F$ of $f$ is the hyperelliptic involution of $F$. Let $\mu:\tilde{S}\rightarrow S$ be the blowing-ups of all the isolated fixed points of $\sigma$ and let $\tilde{\sigma}$ be the induced involution on $\tilde{S}$. The quotient space $\tilde{P}=\tilde{S}/\langle\tilde{\sigma}\rangle$ is a smooth surface, and $f$ induces a ruling $\tilde{h}:\tilde{P}\rightarrow B$ on $\tilde{P}$. Also, the quotient map $\tilde{\pi}:\tilde{S}\rightarrow\tilde{P}$ is a double cover which is determined by the pair $(\tilde{R},\tilde{\delta})$, where $\tilde{R}$ is the branch locus of $\tilde{\pi}$ and $\tilde{\delta}$ is a line bundle such that $\mathcal{O}_{\tilde{P}}(\tilde{R})\cong\tilde{\delta}^{\otimes 2}$ and $\tilde{\pi}_{*}\mathcal{O}_{\tilde{S}}\cong\mathcal{O}_{\tilde{P}}\oplus\tilde{\delta}^{\vee}$. For any contraction $\varphi:\tilde{P}\rightarrow P^{\prime}$ and $R^{\prime}=\varphi(\tilde{R})$, the double cover $\tilde{\pi}$ induces a double cover $S^{\prime}\rightarrow P^{\prime}$, which is determined by $(R^{\prime},\delta^{\prime})$. We call $(R^{\prime},\delta^{\prime})$ the image of $(\tilde{R},\tilde{\delta})$.
@@
We remark that the minimality of $R^{2}$ implies that $s_{g+2}=0$ if $g$ is even (c.f. [LZ17, Thm 2.6 and its proof]) The contraction $\psi$ in Section 2.2 can be decomposed into $\tilde{\psi}:\tilde{P}\rightarrow\hat{P}$ and $\hat{\psi}:\hat{P}\rightarrow P$, where $\tilde{\psi}$ and $\hat{\psi}$ are composed respectively of resolutions of negligible and non-negligible singularities of $R$, cf. [Xia91, Def 4] or [LZ17, Def 2.1]. Let $(\hat{R},\hat{L})$ be the image of $(\tilde{R},\tilde{L})$ in $\hat{P}$. Let $\hat{\psi}=\hat{\psi}_{1}\circ\dotsm\circ\hat{\psi}_{t}$ be the decomposition of $\hat{\psi}$, where $\hat{\psi}_{i}:\hat{P}_{i}\rightarrow\hat{P}_{i-1}$ is a blowing-up at $y_{i-1}$, $\hat{P}_{0}=P$ and $\hat{P}_{t}=\hat{P}$. Let $\hat{R}_{i}$ be the image of $\hat{R}$ in $\hat{P}_{i}$. We remark that the decomposition of $\hat{\psi}$ is not unique. If $y_{i-1}$ is the singular point if $\hat{R}_{i-1}$ of odd multiplicity $2k+1~{}(k\geq 1)$ and there is a unique singular point $y$ of $\hat{R}_{i}$ on the exceptional curve ^$\hat{\mathcal{E}}_{i}$ of multiplicity $22k+$ 2, then we always assume that $\hat{\psi}_{i+1}:\hat{P}_{i+1}\rightarrow\hat{P}_{i}$ is the standard blowing-up at $y_{i}=y$. We call such a pair $(y_{i-1},y_{i})$ a singularity of type $(2k+1\rightarrow 2k+1)$ and call $y_{i-1}$ (resp. $y_{i}$) the first (resp. second) component of such a singularity.
@@
For any singular fiber $F$ of $f$ and $23\leq i\leq g+$ 2, the $i$-th singularity index of $F$ is defined as follows (with respect to the contraction $\psi$): (1) if $i$ is odd, $s_{i}(F)$ equals the number of $(i\rightarrow i)$ type singularities of $R$ over the image $f(F)$; (2) if $i$ is even, $s_{i}(F)$ equals the number of singularities of multiplicity $i$ or $i+1$ of $R$ over the image $f(F)$, neither belonging to the second component of $(i-1\rightarrow i-1)$ type singularities nor to the first component of $(i+1\rightarrow i+1)$ type singularities.
Let $K_{\hat{P}/B}=K_{\hat{P}}-h^{*}K_{B}$ and $R^{\prime}={\hat{R}}\setminus{\hat{V}}$, where $\hat{V}$ is the union of isolated vertical $(-2)$-curves in $\hat{R}$. Here a curve $C\subset\hat{R}$ is said to be isolated in $R$ if there is no other curve $C^{\prime}\subset\hat{R}$ such that $C\cap C^{\prime}\neq\emptyset$. We define
@@
Note that $s_{i}$ is nonnegative for $i\geq 3$ by definition, but it is not clear whether $s_{2}$ is nonnegative or not.
@@
On the other hand we can write $R\sim-(g+1)K_{P/B}+nF$, where $\sim$’ stands for numerical equivalence, and $n=\frac{R^{2}}{4(g+1)}$ is an integer. The following formulas for hyperelliptic fibrations are due to Xiao; we refer to [Xia91, Thm 1] or [LZ17, Thm 2.6] for a proof.
@@
Suppose that $f$ is locally non-trivial. Then the above formulas implies in particular that
@@
Let $f:S\rightarrow B$ be a locally non-trivial hyperelliptic fibration of genus $g\geq 2$. Then the slope $\lambda_{f}$ satisfies
@@
$K_{f}^{2}=K_{S}^{2},\qquad\chi_{f}=\chi=\chi(\mathcal{O}_{S}),\qquad e_{f}=\chi_{top}(S).$ (2.4)

### v3

# Upper bounds on the genus of Hyperelliptic Albanese fibrations

@@ Abstract.
Let $S$ be a minimal irregular surface of general type, whose Albanese map induces a hyperelliptic fibration $f:\,S\to B$ of genus $g$. We prove a quadratic upper bound on the genus $g$, i.e., $g\leq h\big{(}\chi(\mathcal{O}_{S})\big{)}$, where $h$ is a quadratic function. We also construct examples showing that the quadratic upper bounds can not be improved to linear ones. In the special case when $p_{g}(S)=q(S)=1$, we show that $g\leq 14$.
@@
This work is supported by National Natural Science Foundation of China, Shanghai Pilot Program for Basic Research (No. TQ20240202), Fundamental Research Funds for central Universities, Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014) and Natural Science Foundation of Shandong Province (No. ZR2023QA00).
@@ 1. Introduction
We work over the complex number throughout this paper. Let $S$ be a minimal irregular surface of general type, and $a:\,S\to\mathrm{Alb}(S)$ be its Albanese map. We are interested in the case when the image $a(S)$ is a curve. In this case, the Albanese map induces a fibration, which we call the Albanese fibration of $S$:
@@
In fact, by the universal property of the Albanese map, $B\cong a(S)$, and under this isomorphism the above fibration $f$ is nothing but the Albanese map of $S$.
Let $g$ be the genus of a general fiber of $f$. A natural problem is to study the behavior of the genus $g$.
@@
By [Cat00], it is known that the genus $g$ of the Albanese fibration of $S$ is a differential invariant, and hence is also a deformation invariant. Fixing $\chi(\mathcal{O}_{S})$ and $K_{S}^{2}$, there are only finitely many deformation equivalence classes of such surfaces, cf. [BHPV04, § VII]. Hence there must be an upper bound on $g$ depending on $\chi(\mathcal{O}_{S})$ or $K_{S}^{2}$. However, it is still unclear how the upper bound depends on $\chi(\mathcal{O}_{S})$ (or $K_{S}^{2}$). Explicit upper bounds are only known under some extra assumptions.
@@
    Suppose that $g(B)=1$ and $K_{S}^{2}=4\chi(\mathcal{O}_{S})$. We showed that $g\leq\max\{6,3\chi(\mathcal{O}_{S})+1\}$, cf. [LL25]. See also [Ish05] for the case when $f$ is moreover hyperelliptic, where it was proved that $g\leq 2\chi(\mathcal{O}_{S})+2$.
@@
Let $S$ be a minimal irregular surface of general type with $q(S)=1$ and $K_{S}^{2}>4\chi(\mathcal{O}_{S})$, and $f:\,S\to B$ its Albanese fibration whose general fiber is of genus $g$. Suppose that $f$ is hyperelliptic. Then
$g\leq\frac{25}{4}{\chi(\mathcal{O}_{S})}^{2}+\frac{19}{2}\chi(\mathcal{O}_{S})+2.$ (1.1)
We will construct in Section 5 examples of minimal irregular surfaces of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$ admitting hyperelliptic Albanese fibrations, whose general fiber is of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$. This is different from the phenomenons when $K_{S}^{2}\leq 4\chi(\mathcal{O}_{S})$, where the upper bounds are all linear in $\chi(\mathcal{O}_{S})$ as mentioned above.
@@
As mentioned above, the genus $g$ is a deformation invariant when $f$ is the Albanese fibration. Hence one may expect an upper bound on $g$ using invariants of $S$. In general, there is no such upper bound on $g$ if $f$ is not the Albanese fibration. In fact, Penegini-Polizzi ([PP17]) constructed a minimal surface $S$ with $p_{g}(S)=q(S)=2,K_{S}^{2}=4$ (whose Albanese map is generically finite), on which there are fibrations $f_{k}:\,S\to B$, such that the genera of $f_{k}$’s can be arbitrarily large.
Nevertheless, in the case when $f:\,S\to B$ is hyperelliptic, the quadratic upper bound proved in Theorem 1.1 holds true even if $f$ is not the Albanese fibration, and hence our results can be applied in a more general situation. See Theorem 3.1 and Theorem 3.2 for more details.
Another interest of ours is the classification of minimal irregular surface of general type with $\chi(\mathcal{O}_{S})=1$. Although the value $\chi(\mathcal{O}_{S})=1$ is the minimal possibility among surfaces of general type, the classification of such surfaces is widely open. Beauville proved in an appendix to [Deb82] that $p_{g}(S)=q(S)\leq 4$ for such surfaces, and the equality holds if and only if $S$ is isomorphic to a product of two curves of genus $2$.
We are more interested in the case when the Albanese map of $S$ induces a fibration. In this case, $p_{g}(S)=q(S)\leq 3$ by Beauville’s result above. In a series of works [CCL98, HP02, Pen11, Pir02, Zuc03], a full classification has been obtained when $p_{g}(S)=q(S)=3$ or $p_{g}(S)=q(S)=2$. However, the case when $p_{g}(S)=q(S)=1$ seems to be much more mysterious. An explicit upper bound on the genus $g$ of the Albanese fibration is still unknown, although theoretically such an upper bound exists since there are at most finitely many deformation equivalence classes of such surfaces. By Theorem 1.1 above, $g\leq 17$ if the Albanese fibration is hyperelliptic. In fact, we can even get more.
@@
Let $S$ be a minimal surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ of $S$ is hyperelliptic of genus $g$. Then $g\leq 14$. More precisely, it holds $g\leq 10$ except for the following two possible cases:
(i) $K^{2}=8,~{}g=14$;     (ii) $K^{2}=8,~{}g=11$.
The known examples of surfaces with $p_{g}(S)=q(S)=1$ are mostly restricted to the region $g\leq K_{S}^{2}$. The first example with $g>K_{S}^{2}$ was constructed in [FP15], in which $g=7$ and $K_{S}^{2}=6$. More recently, an example with $g=19$ and $K_{S}^{2}=9$ was constructed in [CKY17]. By our result above, the Albanese fibration of the example in [CKY17] must be non-hyperelliptic.
The paper is organized as follows. In Section 2, we mainly review some basic facts about surface fibrations and do some technical preparations. In Section 3 and Section 4, we prove Theorem 1.1 and Theorem 1.2 respectively. Finally in Section 5, we construct examples of irregular surfaces of general type with hyperelliptic fibrations of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$.
Acknowledgments. We are grateful to the referees for valuable comments and suggestions, which make our paper more clear.
@@ 2. Preliminaries
In this section, we mainly review some basic facts and fix the notations. In Section 2.1, we recall some general facts about surface fibrations, and refer to [BHPV04] for more details; and then in Section 2.2 we restrict ourselves to the theory on the hyperelliptic fibrations, which goes back to Horikawa and Xiao, cf. [Hor77, Xia91, Xia92].
@@ 2.1. The surface fibrations
Let $f:S\rightarrow B$ be a fibration of curves of genus $g\geq 2$, i.e. $f$ is a proper morphism from a smooth projective surface $S$ onto a smooth projective curve $B$ with connected fibers, and the general fiber is a smooth projective curve of genus $g$. The fibration $f$ is called relatively minimal if there is no $(-1)$-curve (i.e. a smooth rational curve with self-intersection $-1$) contained in the fibers of $f$. It is called hyperelliptic if its general fiber is hyperelliptic, smooth if all its fibers are smooth, isotrivial if all its smooth fibers are mutually isomorphic, and locally trivial if it is both smooth and isotrivial.
Let $\omega_{S}$ (resp. $K_{S}$) be the canonical sheaf (resp. canonical divisor) of $S$, and let $\omega_{S/B}=\omega_{S}\otimes f^{*}\omega_{B}^{\vee}$ (resp. $K_{f}=K_{S/B}=K_{S}-f^{*}K_{B}$) be the relative canonical sheaf (resp. the relative canonical divisor) of $f$. Put $b:=g(B)$, $p_{g}:=h^{0}(S,\omega_{S})$, $q:=h^{1}(S,\omega_{S})$, $\chi=\chi(\mathcal{O}_{S}):=p_{g}-q+1$, and let $e(S)$ be the topological Euler characteristic of $S$. The basic invariants of $f$ are:
@@
    $e_{f}=\sum e_{F}$, where $e_{F}:=e(F_{red})-(2-2g)$ for any fiber $F$, $F_{red}$ is the reduced part of $F$ and $e(F_{red})$ is the topological Euler characteristic of $F_{red}$. Moreover, $e_{F}\geq 0$, and the equality holds if and only if $F$ is smooth. Hence $e_{f}\geq 0$, and $e_{f}=0$ if and only if $f$ is smooth.
@@ 2.2. Invariants of hyperelliptic fibrations
In this subsection, we restrict to hyperelliptic fibrations. Let $f:S\rightarrow B$ be a relatively minimal hyperelliptic fibration. The relative canonical map of $f$ is generically of degree 2. Hence it determines an involution $\sigma$ on $S$ whose restriction on a general fiber $F$ of $f$ is the hyperelliptic involution of $F$.
Let $\mu:\tilde{S}\rightarrow S$ be the blow-ups of all the isolated fixed points of $\sigma$ and let $\tilde{\sigma}$ be the induced involution on $\tilde{S}$. The quotient space $\tilde{P}=\tilde{S}/\langle\tilde{\sigma}\rangle$ is a smooth surface, and $f$ induces a ruling $\tilde{h}:\tilde{P}\rightarrow B$ on $\tilde{P}$. Also, the quotient map $\tilde{\pi}:\tilde{S}\rightarrow\tilde{P}$ is a double cover which is determined by the pair $(\tilde{R},\tilde{\delta})$, where $\tilde{R}$ is the branch locus of $\tilde{\pi}$ and $\tilde{\delta}$ is a line bundle such that $\mathcal{O}_{\tilde{P}}(\tilde{R})\cong\tilde{\delta}^{\otimes 2}$ and $\tilde{\pi}_{*}\mathcal{O}_{\tilde{S}}\cong\mathcal{O}_{\tilde{P}}\oplus\tilde{\delta}^{\vee}$. (see [BHPV04] Chapter V.22)
For any contraction $\varphi:\tilde{P}\rightarrow P^{\prime}$ of $(-1)$-curves and $R^{\prime}=\varphi(\tilde{R})$, the double cover $\tilde{\pi}$ induces a double cover $S^{\prime}\rightarrow P^{\prime}$, which is determined by $(R^{\prime},\delta^{\prime})$. We call $(R^{\prime},\delta^{\prime})$ the image of $(\tilde{R},\tilde{\delta})$.
@@
The contraction $\psi$ in Section 2.2 can be decomposed into $\tilde{\psi}:\tilde{P}\rightarrow\hat{P}$ and $\hat{\psi}:\hat{P}\rightarrow P$, where $\tilde{\psi}$ and $\hat{\psi}$ are composed respectively of resolutions of negligible and non-negligible singularities of $R$, cf. [Xia91, Def 4] or [LZ17, Def 2.1]. Let $(\hat{R},\hat{L})$ be the image of $(\tilde{R},\tilde{L})$ in $\hat{P}$. Let $\hat{\psi}=\hat{\psi}_{1}\circ\dotsm\circ\hat{\psi}_{t}$ be the decomposition of $\hat{\psi}$, where $\hat{\psi}_{i}:\hat{P}_{i}\rightarrow\hat{P}_{i-1}$ is the blow-up at $y_{i-1}$, $\hat{P}_{0}=P$ and $\hat{P}_{t}=\hat{P}$. Let $\hat{R}_{i}$ be the image of $\hat{R}$ in $\hat{P}_{i}$. We remark that the decomposition of $\hat{\psi}$ is not unique. If $y_{i-1}$ is a singular point of $\hat{R}_{i-1}$ of odd multiplicity $2k+1~{}(k\geq 1)$ and there is a unique singular point $y$ of $\hat{R}_{i}$ on the exceptional curve ^$\hat{\mathcal{E}}_{i}$ of multiplicity $22k+$ 2, then we always assume that $\hat{\psi}_{i+1}:\hat{P}_{i+1}\rightarrow\hat{P}_{i}$ is the standard blow-up at $y_{i}=y$. We call such a pair $(y_{i-1},y_{i})$ a singularity of type $(2k+1\rightarrow 2k+1)$ and call $y_{i-1}$ (resp. $y_{i}$) the first (resp. second) component of such a singularity.
@@
For any singular fiber $F$ of $f$ and $23\leq i\leq g+$ 2, the $i$-th singularity index of $F$ is defined as follows (with respect to the contraction $\psi$):
(1) if $i$ is odd, $s_{i}(F)$ equals the number of $(i\rightarrow i)$ type singularities of $R$ over the image $f(F)$;
(2) if $i$ is even, $s_{i}(F)$ equals the number of singularities of multiplicity $i$ or $i+1$ of $R$ over the image $f(F)$, neither belonging to the second component of $(i-1\rightarrow i-1)$ type singularities nor to the first component of $(i+1\rightarrow i+1)$ type singularities.
Here we give an example to help to understand the above singularity indices.
[…]
Suppose that $F$ is a singular fiber of $f$, and $\Gamma$ is its corresponding fiber in the geometrical ruled surface $h:\,P\to B$. Let $t$ be a local coordinate of $B$ around $b=h(\Gamma)$, and $x$ be an affine fiber-coordinate of the $\mathbb{P}^{1}$-bundle over $B$.
  1. (1)
    Suppose that the local equation of $R$ over $b$ is $(t^{2k}-x^{2k})(x^{2g+2-2k}-1)=0$ with $2\leq k\leq\big{[}\frac{g+1}{2}\big{]}$. Then $p$ is the unique singularity of $R$ on $\Gamma$ of multiplicity $2k$. In this case, $s_{2k}(F)=1$ and $s_{i}(F)=0$ for all other $i\geq 3$.
  2. (2)
    Suppose that the local equation of $R$ over $b$ is $(t^{2(2k+1)}-x^{2k+1})(x^{2g+1-2k}-1)=0$ with $1\leq k\leq\big{[}\frac{g}{2}\big{]}$. Then $p$ is a singularity of $R$ on $\Gamma$ contained in a singularity of type $(2k+1\to 2k+1)$. In this case, $s_{2k+1}(F)=1$ and $s_{i}(F)=0$ for all other $i\geq 3$.
  3. (3)
    Suppose that $g$ is odd, and that the local equation of $R$ over $b$ is $t(t^{g+1}-x^{2(g+1)})=0$. Then $p$ is also the unique singularity of $R$ on $\Gamma$. In this case, $s_{g+2}(F)=1$ and $s_{i}(F)=0$ for all other $i\geq 3$.
We remark that the minimality of $R^{2}$ implies that $s_{g+2}(F)=0$ if $g$ is even (c.f. [LZ17, Thm 2.6 and its proof])
Let $K_{\hat{P}/B}=K_{\hat{P}}-\hat{h}^{*}K_{B}$ and $R^{\prime}={\hat{R}}\setminus{\hat{V}}$, where $\hat{V}$ is the union of isolated vertical $(-2)$-curves in $\hat{R}$. Here an irreducible curve $C\subset\hat{R}$ is said to be isolated in $R$ if there is no other irreducible curve $C^{\prime}\subset\hat{R}$ such that $C\cap C^{\prime}\neq\emptyset$. We define
@@
Note that $s_{i}$ is nonnegative for $i\geq 3$ by definition, but the singularity index $s_{2}$ might be negative, cf. [GGL25].
@@
According to [Xia91, page 604], we can write $R\sim-(g+1)K_{P/B}+nF$ where $\sim$’ stands for numerical equivalence, and
$n=\frac{R^{2}}{4(g+1)},$ (2.2)
which is an integer. The following formulas for hyperelliptic fibrations are due to Xiao; we refer to [Xia91, Thm 1] or [LZ17, Thm 2.6] for a proof.
@@
Suppose that $f$ is not locally trivial. Then the above formulas imply in particular that
@@
Let $f:S\rightarrow B$ be a not locally trivial hyperelliptic fibration of genus $g\geq 2$. Then the slope $\lambda_{f}$ satisfies
@@
$K_{f}^{2}=K_{S}^{2},\qquad\chi_{f}=\chi=\chi(\mathcal{O}_{S}),\qquad e_{f}=e(S).$ (2.5)